OpenStudy (anonymous):

A spherical balloon with radius r inches has volume V(r) = 4 3 πr3. Find an expression that represents the amount of air required to inflate the balloon from a radius of r inches to a radius of r + 4 inches. (Express your answer in terms of π and r.)

OpenStudy (campbell_st):

for the larger radius \[V(r + 4) = 4/3 \pi (r + 4)^3 \] the smaller radius \[V(r) = 4/3 \pi r^3\] the increase in volume when the radius increases from r to r + 4 is larger radius volume - smaller radius volume \[V = 4/3 \pi[(r + 4)^3 - r^3]\] just expand and simplify

OpenStudy (anonymous):

so the expanded version would be V=4/3pi(r^3+64)-r^3 ?

OpenStudy (anonymous):

Your expansion of (r+4)^3 isn't correct. Multiply it out step by step. Start with (r+4)(r+4)(r+4) and multiply step by step.

OpenStudy (anonymous):

so i got r^3+12r^2+48r+64. is that right? what do i do now?

OpenStudy (anonymous):

Yup. Put that in, gather the like terms, and you will have your answer.

OpenStudy (anonymous):

I think you might get something similar to\[V=\frac{4 \pi}{3}(12r^2+48r+64)=\frac{16 \pi}{3}(3r^2+16r+16)\]

OpenStudy (anonymous):

OOOPS!! Bad factoring on my previous answer! Correction: \[V=\frac{4 \pi}{3}(12r^2+48r+64)=\frac{16 \pi}{3}(3r^2+12r+16)\]Sorry for the careless goof. Do math every day.

OpenStudy (anonymous):

that's the answer i got! thank you so much for your help! (:

OpenStudy (anonymous):

No sweat. Do math every day.